POY11 does not go out of the vacuum
It does but slighty low and does not get on mirrors. We need to change optic mounts to adjust the height. Red is flashing in yarm at 00 and 10 modes. TRY is ~0.4-0.5.
I've adjusted BS angle, camera and TRX PD at ETMX table so I can see red flashing at 03 mode while green is locked to 00 and its transmission is maximized. I thought that by adjusting BS angle, I will be able to align red to 00 not disturbing green, but this was not the case. Maximum TRX I could get was 0.1. I've adjusted POX to get into PD and I can see PDH signal though I can't lock as cavity is still misaligned for red.
We put the POY beam onto the POY PD. The Yarm is currently locked on IR with ~0.65 transmission.